Your brother has offered to give you either $5,000 today or $10,000 in ten years. If the interest rate is 7% per year, which option is preferable?
Answer #1

We can compare two financial alternatives only when we have present or future worth of each alternative. We cannot compare present worth of one alternative with the future worth of another alternative. We can either choose to calculate present or future worth of each of the alternatives. I would prefer to calculate present worth of each case.   

Case 1

Present worth of alternative =PW1= $5000

Case 2

Rate of interest=i=7% per year

Time period=n=10 years

Future Value of offering=FV=$10000

PW of alternative=FV/(1+i)^n=10000/(1+7%)^10=$5083.49

Present worth of second alternative is more. So, I would prefer second alternative i.e. $10000 in ten years.